# arXiv:2404.05610v1[cs.DC] 08 Apr 2024

### 08 Apr 2024 v1

@@ IV-B Graph Algorithms
Currently, most state of the art HPC platforms are mainly designed for numerical applications with fairly regular data access and communication patterns. However, data intensive irregular workloads become more and more frequent, for example, in materials science or data analysis tasks such as (human) brain analysis. Graphs are commonly used to represent data sets in these applications and we therefore require efficient distributed graph algorithms either for network analysis or as important building blocks for more complex applications [25, 26, 27]. To demonstrate the applicability of KaMPIng in this setting, we provide a simple distributed breadth-first search (BFS) implementation in Figure 7. We assume the graph to be distributed among the ranks with each rank holding a subset of the vertices and their incident edges. Locally, the graph is represented as an adjacency array. For each vertex $v$, the bfs returns the distance (number of hops) between the source vertex s and $v$. KaMPIng’s utility function with_flattened(...), which flattens a container of nested destination-message pairs by transforming it into a contiguous range while also providing send counts, proves to be especially useful.
As for the sample sort example, we implemented the distributed BFS algorithm using all previously discussed (C++) MPI bindings in a comparable way. The implementations only differ for the frontier exchange and completion logic, which can be implemented in KaMPIng using only $22$ lines of code, whereas plain MPI requires $46$ lines. Our closest competitor regarding code length is RWTH-MPI with $32$ LOC (see Footnote 6 for all other bindings)  [^7] The code considered here is structured slightly differently compared to Figure 7, which has been shortened for readability. See https://github.com/kamping-site/kamping-examples/tree/main/include/bfs/ for full implementations..
In Figure 8, we provide a performance evaluation of the implementations on a variety of different graph families, which is discussed in more detail in Section V-A.
@@
Figure 8: Running time of BFS using different all-to-all variants.
Graph Partitioning. As a more complex showcase, we integrated KaMPIng into the state-of-the-art distributed multilevel graph partitioner dKaMinPar [28], consisting of roughly 30 000 LOC and including its own abstraction layer with specialized graph-specific communication primitives over plain MPI. The partitioner uses size-constrained label propagation to iteratively cluster and contract the input graph, shrinking it down until its size falls below a certain threshold. Due to the project’s size, we only focus on this component and compare an implementation based on dKaMinPar’s application-specific MPI abstraction layer, a plain MPI-based implementation and a KaMPIng-based implementation. We have extracted the shared code of all implementations (202 LOC) to a base class and only focus on the MPI-heavy part of the algorithm’s implementation. Here, the plain MPI-based implementation (154 LOC) is roughly 17.5% larger than the KaMPIng-based implementation (127 LOC), which in turn is 16.5% larger than the implementation based on dKaMinPar’s specialized abstraction layer (106 LOC). We observed the same running times for all variants.
@@ IV-C Integrating KaMPIng in RAxML-NG
As our largest application benchmark, we consider RAxML-NG. RAxML [29] and its modern rewrite RAxML-NG [10] are widely used real-world programms for phylogenetic inference in the field of bioinformatics with over $50\,000$ citations. RAxML-NG is written in C++and uses a custom non-trivial abstraction layer over pthreads + MPI parallelism with over 700 lines of code. We use KaMPIng to substantially simplify the MPI part of this abstraction layer; demonstrating that KaMPIng can easily be integrated even in large and well-established scientific codes which use hybrid parallelization (for an example, see Figure 9). If KaMPIng would have been available at the time, the RAxML-NG developers would have never have needed to write, unit-test, maintain, and document over hundred lines of complex code.
@@ V-A Sparse and Low-Latency All-To-All communication
All-to-all exchanges are one of the most frequent communication patterns in distributed computing. However, there is a large algorithmic design space for all-to-all communication ranging from algorithms with near optimal communication volume but latency at least linear in the number of processing elements to algorithms following a hypercube communication scheme with logarithmic latency but also a communication volume that is increased by a logarithmic factor [20].
@@
With KaMPIng’s GridCommunicator plugin we go part of the way of trading communication volume for reduced latency by resorting to two-dimensional grid communication [30]. The processors are organized in a virtual two-dimensional grid and messages are routed in two hops to their destination to achieve a message start-up latency in $\mathcal{O}(\sqrt{p})$, where $p$ denotes the size of the communicator. This enables hardware agnostic latency reduction with asymptotic guarantees, in contrast to the variants provided by most MPI implementations.
@@
Additionally, MPI’s standard collectives have not been designed with sparse communication patterns in mind. MPI_Alltoallv, for example, requires a send counts parameter consisting of an array with one entry for each rank of the communicator. This yields a time complexity which is at least linear in the size of the underlying communicator. To mitigate this problem for static communication patterns, neighborhood collectives have been added to MPI-3.0 allowing the user to perform MPI_Alltoall(v) and MPI_Allgather(v) on a previously defined (sparse) graph topology. However, for applications and algorithms with rapidly changing communication partners, e.g., (dynamic) graph algorithms, the overhead of defining a new communication graph topology prior to every few all-to-all exchanges may impose too much overhead. KaMPIng’s SparseAlltoall plugin offers a lightweight alternative for these scenarios. It introduces a sparse_send_buf named parameter accepting any C++container consisting of destination-message pairs, like for example std::unordered_map<int, std::vector<T>>. For the actual data exchange, the plugin uses the NBX algorithm for sparse all-to-all communication by Hoefler et al. [31].
@@
Figure 8 shows an evaluation of our different all-to-all strategies using the weak-scaling BFS benchmark introduced earlier on three different graph families [32], where each rank holds $2^{12}$ vertices and $2^{15}$ edges. Erdős-Rényi graphs possess almost no locality (most edges cross rank boundaries) but a small diameter, whereas random geometric graphs (RGG) are highly local with a high diameter. Regarding locality, random hyperbolic graphs (RHG) range somewhere in between and also have small diameter. In contrast to Erdős-Rényi graphs and RGGs they possess high-degree vertices. In the experiment, we compare different algorithms for the actual frontier exchange in each BFS step: built-in MPI_Alltoallv (mpi, KaMPIng), KaMPIng’s sparse all-to-all plugin, KaMPIng’s grid all-to-all plugin, and MPI_Neighbor_alltoallv (mpi_neighbor). For RHGs (and less pronounced for Erdős-Rényi graphs) the most scalable communication method is our grid all-to-all approach. Grid all-to-all also outperforms built-in MPI_Alltoallv on RGGs. Due to their high diameter and locality, a competitive performance on RGGs can only be achieved with sparse communication. KaMPIng’s sparse all-to-all approach is only slightly slower than MPI_Neighbor_alltoallv. Note that when rebuilding MPI’s communication graph before each frontier exchange, which simulates dynamic communication patterns to some extent, MPI_Neighbor_alltoallv does not scale.
@@ V-B User-Level Failure Mitigation
With the increasing number of processors in high performance computing clusters, the probability that some processors fail during a computation rises. Handling such failures constitutes a major challenge for future exascale systems [33]. In upcoming systems, we expect a hardware failure to occur every 30 to 60 minutes [34, 35, 36]. The upcoming MPI 5.0 standard enables developers to develop software able to recover from such failures by employing User-Level Failure-Mitigation (ULFM) [37]. As one example of the potency of our plugin architecture (Section III-F) we developed an abstraction layer over ULFM supporting all functions of the proposal [^9] https://fault-tolerance.org/. As plugins can add custom error handling hooks, this enables users to develop fault-tolerant algorithms using idiomatic C++exceptions instead of checking return codes.
@@ VI Conclusion
KaMPIngis open source, extensively tested, and currently used in multiple research projects. It covers MPI’s collective and (nonblocking) point-to-point operations, the most commonly used MPI features [2]. In the future, we plan to extend the standard coverage while also working towards our goal of building a basic algorithmic toolbox on top of it, to ease rapid prototyping and analysis of distributed algorithms with a strong focus on performance.
In future releases of KaMPIng, we plan to generalize our all-to-all primitives to higher dimensions and integrate mechanism for message aggregation for request-reply patterns when reading from globally distributed data. With distributed containers, we want to enable lightweight bulk parallel computation inspired by MapReduce [40] and Thrill [41], while not locking the programmer into the walled garden of a particular framework. We strive to establish KaMPIng as a stable core for a whole ecosystem of future general purpose distributed algorithms and applications.

### 23 Aug 2024 v2

@@ IV-B Graph Algorithms
Currently, most state-of-the-art HPC platforms are mainly designed for numerical applications with fairly regular data access and communication patterns. However, data-intensive irregular workloads become more and more frequent, for example, in materials science or data analysis tasks such as (human) brain analysis. Graphs are commonly used to represent data sets in these applications and we therefore require efficient distributed graph algorithms either for network analysis or as important building blocks for more complex applications [28, 29, 30]. To demonstrate the applicability of KaMPIng in this setting, we provide a simple distributed breadth-first search (BFS) implementation in Figure 9. We assume the graph to be distributed among the ranks with each rank holding a subset of the vertices and their incident edges. Locally, the graph is represented as an adjacency array. For each vertex $v$, the bfs returns the distance (number of hops) between the source vertex s and $v$. KaMPIng’s utility function with_flattened(...), which flattens a container of nested destination-message pairs by transforming it into a contiguous range while also providing send counts, proves to be especially useful.
As for the sample sort example, we implemented the distributed BFS algorithm using all previously discussed (C++) MPI bindings comparably. The implementations only differ for the frontier exchange and completion logic, which can be implemented in KaMPIng using only $22$ lines of code, whereas plain MPI requires $46$ lines. Our closest competitor regarding code length is RWTH-MPI with $32$ LOC (see Footnote 7 for all other bindings)  [^8] The code considered here is structured slightly differently compared to Figure 9, which has been shortened for readability. See https://github.com/kamping-site/kamping-examples/tree/main/include/bfs/ for full implementations..
In Figure 10 we evaluate these implementations using a variety of different graph families and observe that KaMPIng introduces no additional overhead compared to plain MPI. KaMPIng also provides optimized collective operations, which provide better scalability than MPI_Alltoallv on some graph families, which we discuss in Section V-A.
@@
Figure 10: Running time of BFS comparing MPI and KaMPIng and additional all-to-all implementations.
Graph Partitioning
As a more complex showcase, we integrated KaMPIng into the state-of-the-art distributed multilevel graph partitioner dKaMinPar [31], consisting of roughly 30 000 LOC and including its own abstraction layer with specialized graph-specific communication primitives over plain MPI. The partitioner uses size-constrained label propagation to iteratively cluster and contract the input graph, shrinking it down until its size falls below a certain threshold. Due to the project’s size, we only focus on this component and compare an implementation based on dKaMinPar’s application-specific MPI abstraction layer, a plain MPI-based implementation and a KaMPIng-based implementation. We have extracted the shared code of all implementations (202 LOC) to a base class and only focus on the MPI-heavy part of the algorithm’s implementation. Here, the plain MPI-based implementation (154 LOC) is roughly 17.5% larger than the KaMPIng-based implementation (127 LOC), which in turn is 16.5% larger than the implementation based on dKaMinPar’s specialized abstraction layer (106 LOC). We observed the same running times for all variants.
@@ IV-C Integrating KaMPIng in RAxML-NG
As our largest application benchmark, we consider RAxML-NG. RAxML [32] and its modern rewrite RAxML-NG [13] are widely used real-world programs for phylogenetic inference in the field of bioinformatics with over $50\,000$ citations. RAxML-NG is written in C++ and uses a custom non-trivial abstraction layer over pthreads + MPI parallelism with over 700 lines of code. We use KaMPIng to substantially simplify the MPI part of this abstraction layer; demonstrating that KaMPIng can easily be integrated even in large and well-established scientific codes which use hybrid parallelization (for an example, see Figure 11). If KaMPIng had been available at the time, the RAxML-NG developers would have never needed to write, unit-test, maintain, and document over a hundred lines of complex code [^9] Here, MPI and application code are heavily intertwined, making a fair comparison for exact LOC hard..
@@ V-A Sparse and Low-Latency All-To-All communication
All-to-all exchanges are one of the most frequent communication patterns in distributed computing. However, there is a large algorithmic design space for all-to-all communication ranging from algorithms with near-optimal communication volume but latency at least linear in the number of processing elements to algorithms following a hypercube communication scheme with logarithmic latency but also a communication volume that is increased by a logarithmic factor [23].
@@
With KaMPIng’s GridCommunicator plugin we go part of the way of trading communication volume for reduced latency by resorting to two-dimensional grid communication [33]. The processors are organized in a virtual two-dimensional grid and messages are routed in two hops to their destination to achieve a message start-up latency in $\mathcal{O}(\sqrt{p})$, where $p$ denotes the size of the communicator. This enables hardware-agnostic latency reduction with asymptotic guarantees, in contrast to the variants provided by most MPI implementations.
@@
Additionally, MPI’s standard collectives have not been designed with sparse communication patterns in mind. MPI_Alltoallv, for example, requires a send counts parameter consisting of an array with one entry for each rank of the communicator, yielding a time complexity linear in the communicator size. To mitigate this problem for static communication patterns, neighborhood collectives have been added to MPI-3.0 allowing the user to perform MPI_Alltoall(v) and MPI_Allgather(v) on a previously defined (sparse) graph topology. However, for applications and algorithms with rapidly changing communication partners, e.g., (dynamic) graph algorithms, the overhead of defining a new communication graph topology every few all-to-all exchanges may impose too much overhead. KaMPIng’s SparseAlltoall plugin offers a lightweight alternative for these scenarios, accepting a set of destination-message pairs as argument. For the actual data exchange, the plugin uses the NBX algorithm for sparse all-to-all communication by Hoefler et al. [34].
@@
Both techniques for improving irregular sparse exchanges work especially well for improving scalability of distributed graph algorithms [35, 36]. Figure 10 shows an evaluation of our different all-to-all strategies using the weak-scaling BFS benchmark introduced earlier on three different graph families [37], where each rank holds $2^{12}$ vertices and $2^{15}$ edges. Erdős-Rényi graphs possess almost no locality (most edges cross rank boundaries) but small diameter, whereas random geometric graphs (RGG) are highly local with a high diameter. Regarding locality, random hyperbolic graphs (RHG) range somewhere in between and also have small diameter. In contrast to Erdős-Rényi graphs and RGGs they possess high-degree vertices. In the experiment, we compare different algorithms for the actual frontier exchange in each BFS step: built-in MPI_Alltoallv (mpi, KaMPIng), KaMPIng’s sparse all-to-all plugin, KaMPIng’s grid all-to-all plugin, and MPI_Neighbor_alltoallv (mpi_neighbor). For RHGs (and less pronounced for Erdős-Rényi graphs) the most scalable communication method is our grid all-to-all approach. Grid all-to-all also outperforms built-in MPI_Alltoallv on RGGs. Due to their high diameter and locality, a competitive performance on RGGs can only be achieved with sparse communication. KaMPIng’s sparse all-to-all approach is only slightly slower than MPI_Neighbor_alltoallv. Note that when rebuilding MPI’s communication graph before each frontier exchange, which simulates dynamic communication patterns to some extent, MPI_Neighbor_alltoallv does not scale.
@@ V-B User-Level Failure Mitigation
With the increasing number of processors in high-performance computing clusters, the probability that some processors fail during a computation rises. Handling such failures constitutes a major challenge for future exascale systems [38]. In upcoming systems, we expect a hardware failure to occur every 30 to 60 minutes [39, 40, 41]. The upcoming MPI 5.0 standard enables developers to develop software able to recover from such failures by employing User-Level Failure-Mitigation (ULFM) [42]. As one example of the potency of our plugin architecture (Section III-F) we developed an abstraction layer over ULFM supporting all functions of the proposal [^11] https://fault-tolerance.org/. As plugins can add custom error handling hooks, this enables users to develop fault-tolerant algorithms using idiomatic C++ exceptions instead of checking return codes, as shown in Figure 12.
try {
comm.allreduce(/* ... */);
} catch ([[maybe_unused]] MPIFailureDetected const& _) {
if (!comm.is_revoked()) {
comm.revoke();
// Create a new communicator containing only the
// surviving processes.
comm = comm.shrink();
Figure 12: Handling a process failure using KaMPIng’s ULFM plugin.
@@ VI Conclusion
KaMPIngis open source, extensively tested, and currently used in multiple research projects. In the future, we plan to extend the standard coverage while also working towards our goal of building a basic algorithmic toolbox on top of it, to ease rapid prototyping and analysis of distributed algorithms with a strong focus on performance. To this end, we will also further elicit the needs of the MPI community.
We are currently working on generalizing the indirection patterns for all-to-all primitives to higher dimensions, while also incorporating message aggregation. This is applicable in both request-replay patterns when reading from globally distributed data, and algorithms with highly-irregular communication without hard synchronization. We are implementing these building blocks on top of KaMPIng and will integrate them into future library releases.
With distributed containers, we want to enable lightweight bulk parallel computation inspired by MapReduce [45] and Thrill [46], while not locking the programmer into the walled garden of a particular framework. We strive to establish KaMPIng as a stable core for a whole ecosystem of future general-purpose distributed algorithms and applications.
@@ Acknowledgment
This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research.
ChatGPT-3 was used to verify the grammatical correctness of a handful of sentences. No generated text has been used in the manuscript. GitHub Copilot was used to assist in writing unit tests and code documentation, and for minor code-completion. No code generation has been used for interface, library, and algorithm design.
We thank Marvin Williams for helping us with his C++ expertise and assistance in C++ standard-related matters.